OpenStudy (anonymous):

1(12) = 4200[1+7/1]^(1*12)

OpenStudy (anonymous):

1*12 = 4200(1+7/1)^(1*12) 12=4200(8)^(12) calculate it through calculator

OpenStudy (anonymous):

ok thx

OpenStudy (a_clan):

LHS is to be found . Its the final amount

OpenStudy (anonymous):

it wont calculate

OpenStudy (anonymous):

A_clan can u help me

OpenStudy (a_clan):

It is A=4200[1+ 7/100]^(12)

OpenStudy (anonymous):

ok thx i wrote it wrong

OpenStudy (a_clan):

seems you misunderstood the formula

OpenStudy (anonymous):

sorry